How do I calculate Financial Leverage given these numbers?

(amounts in millions) 2015 2014
Balance sheet
Total assets $96,296 $94,809
Shareholders' equity 6,647 9,040
Income statement
Net sales $102,780 $97,428
Net income 5,426 5,696
Answer #1

Financial leverage formula = Average total assets/Average equity

Average total assets = (96296+94809/2) = $95552.50

Average equity = (6647+9040/2) = $7843.50

Financial leverage = 95552.50/7843.50 = 12.18 Times